# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v1

@@ Abstract
This paper describes a novel methodology for determining representational alignment, developed upon the existing Spotlight Resonance method. Using this, it is found that algebraic symmetries of network primitives are a strong predictor for task-agnostic structure in representations. Particularly, this new tool is used to gain insight into how discrete representations can form and arrange in autoencoder models, through an ablation study where only the activation function is altered. Representations are found to tend to discretise when the activation functions are defined through a discrete algebraic permutation-equivariant symmetry. In contrast, they remain continuous under a continuous algebraic orthogonal-equivariant definition. These findings corroborate the hypothesis that functional form choices can carry unintended inductive biases which produce task-independent artefactual structures in representations, particularly that contemporary forms induce discretisation of otherwise continuous structure — a quantisation effect. Moreover, this supports a general causal model for one mode in which discrete representations may form, and could constitute a prerequisite for downstream interpretability phenomena, including grandmother neurons, discrete coding schemes, general linear features and possibly Superposition. Hence, this tool and proposed mechanism for the influence of functional form on representations may provide several insights into emergent interpretability research. Finally, preliminary results indicate that quantisation of representations appears to correlate with a measurable increase in reconstruction error, reinforcing previous conjectures that this collapse can be detrimental.
@@ 1.1 How Functional Form Choices May Result in Representational Inductive Biases
Many of the primitive functions in contemporary deep learning produce maps that differ as the direction changes; a property termed as anisotropy. This angular variation of the functions is hypothesised to make specific directions distinct, which the network then alters its representations about through optimisation. Hence, an angularly varying map could be expected to induce shaped angular representations. It is this connection between functional forms and representations which shall be explored in this work.
Directions singled out by a functional form’s structure may be termed a ‘privileged basis’ [1] or a set of ‘distinguished directions’ [2]. Such distinguished directions are commonplace through activation functions, normalisers, initialisers, regularisers, optimisers, architectures, operations, and more [2]. Hence, there are several avenues hypothesised through which representational influence may occur: directly through functional forms, indirectly through optimiser forms, and inherently through neuron connectivities. This proposed triad of influences would suggest that the representational alignment phenomenon may be systematic to deep learning’s anisotropic primitives. As a starting point, this work will isolate and determine the direct influence of activation functions as one of these causal influences on representations.
In particular, contemporary functional forms tend to make distinct the standard basis’ vectors, $\hat{e}_{i}$, and in such a way that each basis vector’s distinction is equal. This particular form of anisotropy can be categorised through their shared, underlying permutation symmetry construction [2]. This can act as a defining characteristic relation for most contemporary functional forms. This is often due to their elementwise application, but appears more generally. The specific relation depends on the primitive in question; for most activation functions, this is an algebraic permutation equivariance [3, 2], as displayed in Eqn. 1. The hypothesis is deepened by suggesting that the maximal symmetries to which functional forms belong are a strong predictor of representational structure.
@@
The Spotlight Resonance Method [4] has already empirically demonstrated how such activation functions can produce a tendency for representational alignments about the distinguished directions. This was achieved through transforms, and altering the number of the functional form’s distinguished directions, and it observed corresponding changes in the representations mirroring the transformed structure of the distinguished directions. The conclusion was that functional forms stimulate this tendency towards representational alignment, producing emergent task-agnostic alignment structure in representations, and hence establishing a causal relationship between the two. The observed structure consisted of a tendency for a greater density of representations to cluster about the functional form’s distinguished basis. This work also indicated the presence of ’grandmother neurons’, indicating how many clusters represent a distinct semantic and also associated their alignment with functional form choices.
@@ 1.2 Motivations for This Study
Isotropic Deep Learning [2] takes the consequences of such observations a step further. The Spotlight Resonance Method determined that task-agnostic structure in representations is associated with a functional form’s distinguished directions. The Isotropic Deep Learning approach proposes that since this structure transforms with the functional form, the anisotropic functional form choices may be the fundamental causal mode that induces the array of interpretable structures to begin with. In effect, anisotropic primitives provide an absolute frame, for each representation space, about which the representations organise and discretise into distinct clusters. Hence, questioning whether the ubiquity of anisotropy in contemporary deep learning constitutes an underlying origin for many such interpretable and emergent-structure phenomena, which are commonly observed — and whether this is beneficial or not. Furthermore, determining the validity of such a hypothesis would indicate whether these phenomena are truly fundamental to deep learning or stimulated by, and contingent on, the specific choices in functional form. This could have significant implications for the approach of network interpretability.
This position motivates the comparative study of this work. Determining whether one can control and demonstrate the emergence and elimination of structure through broadening deep learning’s primitives to other form-defining choices, each providing its own inductive bias. One such proposed choice [2] is by redefining primitives through an orthogonal algebraic-equivariance, displayed in Eqn. 2. This is termed an ‘isotropic’ definition of forms, and is a supergroup-symmetry of the original permutation symmetry, $S_{n}\subset\mathrm{SO}\left(n\right)$. This defining choice of isotropic primitives is interesting, since it can eliminate all form-distinguished directions in a network. Therefore, functions developed to adhere to this modified symmetry definition can be used as a comparative mode to determine whether functional forms can induce these interpretable structures and indicate whether downstream alignment phenomena are attributable to anisotropic choices.
$\forall\mathbf{R}\in\mathrm{SO}\left(n\right),\forall\vec{x}\in\mathbb{R}^{n}\quad\mathbf{f}\left(\mathbf{R}\vec{x}\right)=\mathbf{R}\mathbf{f}\left(\vec{x}\right)$ (2)
@@
In particular, anisotropic forms may lead to a tendency for representations to overly cluster around distinguished directions. This is due to the anisotropic functional form’s nature, where some angular arrangements may become overrepresented through optimisation due to their irregularity. This task-agnostic over-density bias may be detrimental, as representations may tend to collapse towards single directions, such as aligned or anti-aligned, and consequently lose informative representational degrees of freedom, potentially damaging a network’s internal expressiveness of otherwise continuous structure. Hence, anisotropy may result in a quantising effect on representations, and thus also the semantics being represented. Consequently, isotropic primitives are expected to remove this form-induced task-agnostic discretising bias on representations.
The pathological aspect arises from the hypothesis that the underlying distributions might be better represented continuously, as this aligns with the frequently continuous real-world semantics being represented, such as colour gradients, lighting shades, positions and poses of objects, and broad object morphologies. Therefore, enabling latent variables and representations to be continuous and potentially more interpolatable at various levels through a network, instead of an imposed task-agnostic structure, may be beneficial for many tasks. This may increase representational capacity and organisation of the representations.
@@
This connects to the coding scheme of the network. Axis-aligned representations, previously observed to frequently emerge [4], approximate a local coding scheme for the network. This is where one neuron corresponds to one semantic. This coding scheme has decreased concept-interference, but has a much lower representational capacity than other coding schemes. Models have been demonstrated to produce other interference-capacity balances with various input sparsities [1, 4], suggesting that axis-alignment is not an appropriate universal inductive bias — further motivating the study of alternative functional form inductive biases and their effects on representations.
@@
A functional form that does not distinguish particular directions may therefore not deter such continuous and interpolatable structures present in the dataset, and their preservation may be advantageous to the network. Yet, sometimes discretisation may be appropriate, such as in classification networks, where abstracting away redundant detail can be beneficial — though in autoencoding models this would not be an expectation. In any case, rigidly imposing this through a task-agnostic inductive bias about a specific arrangement might be detrimental. For example, when reducing the dimensionality of a dataset inherently containing discrete representations via a linear layer, to some arbitrary lower dimension, one would not expect it always appropriate for those discrete directions to be best represented through an alignment to an orthonormal basis or other imposed geometry for that space — as these are as arbitrary to the bottleneck chosen.
@@
Other speculative pathological modes are attributed to functional form quantisation. For example, discrete representations may be particularly susceptible to perturbations which move activations outside of the cluster, into an in-between region of representation space. Maps from these unpopulated regions may be poorly trained and fragile, due to their training-time rarity, and therefore highly unpredictable in their map. This may suggest that such a quantising inductive bias decreases adversarial robustness, whereas removing this bias through isotropy may be advantageous. However, this speculation is not explored within this work.
This paper aims to assess the validity of the former prediction. The motivating questions are: do discrete-symmetry defined functional forms tend to produce approximately discrete representations, quantising an otherwise underlying continuous data structure, whilst continuous-symmetry defined functional forms may tend to preserve a more continuous representation. Furthermore, is the prior construction potentially pathological in nature? Do continuous representations tend to improve performance, since quantised representations may harm expressiveness? Therefore, determining the validity of this prediction may provide crucial insight into the potential of Isotropic deep learning and the emergence of downstream interpretability structure, namely alignment, grandmother neurons, general interpretable directions, discrete codes, and possibly Superposition [1] and Neural Collapse [5].
@@
Most importantly, this study aims to clarify the link between functional form algebraic symmetries and their impact on representations. It will first be demonstrated that symmetry is a key predictor of the emergent task-agnostic structure in representations. This encourages an exploration into the effects of various symmetry groups and possibly a reimagining of functional form primitives, network-wide, if alternative symmetries are beneficial. This is because symmetry can act as a classifying and definitional approach to functional forms.
@@ 1.3 Methodological Approach
Instead of rotating the privileged basis or altering the number of distinguished directions to determine alignment of representations, this study will compare representations between the presence and absence of any distinguished directions to determine if this is a causal mechanism for the production of approximately discrete representations. Due to the expected systematic nature of this inductive bias, care was taken to prevent the unintentional reintroduction of anisotropy. To achieve this, distinguished directions will be removed network-wide to avoid these confounding and obscuring interactions. This is achieved by utilising isotropic primitives and an autoencoder model for reconstruction  [^1] The use of an autoencoder model prevents anisotropy through choices in the output layer representation. For example, one-hot encoding in classification outputs is also a choice in form defined through permutation symmetry; therefore, it could induce further spurious structure. A reconstruction task using autoencoders may therefore mitigate this and enable more data-representative and ‘natural’ embeddings. Moreover, it has been speculated [4, 2] whether the neural-collapse classification phenomenon [5] is related to the proposed mechanisms. So autoencoder reconstruction provided a suitable minimalistic setting, free from these influences, to study form-induced representational quantisation., except for an activation function which can selectively reintroduce anisotropy through changes to its form. Hence, this isolates the causality of the phenomenon to the activation functions.
This requires two activation functions, which are highly similar in action but differ in the presence or absence of anisotropy. The isotropic position paper [2] provides such a preliminary isotropic activation function, which enables this comparative analysis. An analogue of standard elementwise (anisotropic) tanh, displayed in Eqn. 3, can be compared with its orthogonal analogue Isotropic-Tanh, displayed in Eqn. 4 — each is displayed in multivariate form for comparison $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. These functions produce equivalent non-linear maps along the standard basis directions: $f_{\text{iso}}\left(\alpha\hat{e}_{i}\right)=f_{\text{aniso}}\left(\alpha\hat{e}_{i}\right)$ forall $\hat{e}_{i}$ and $\alpha$. Standard-Tanh is algebraically equivariant to the hyperoctahedral group (standard-basis permutation with sign-flips), $B_{n}$, whilst isotropic-tanh is algebraically equivariant to the orthogonal group, $\mathrm{O}\left(n\right)$, with $B_{n}\subset\mathrm{O}\left(n\right)$.
@@
This approach is also extended to a Leaky-ReLU [6] family, which are modified to embody the following symmetries $S_{n}\subset D_{n}\subset\mathrm{O}\left(n\right)$. This is discussed further in App. B.4 and indicates that algebraic symmetries are the primary predictor, since all other algebraic properties of Leaky-ReLU substantially differ from the tanh-like family of Eqns. 3 and 4.
This work also modifies the prior Spotlight Resonance Method (SRM). This modified tool is termed the Privileged-Plane Projective Method (PPP method), which operates on a similar principle. However, unlike SRM, this adaptation does not project out the vector length in the process. Hence, information regarding both angular and magnitude distributions is retained. This provides a more nuanced understanding of the network’s representational structures; however, SRM may still be preferable in determining statistical measures of the angular representations. Furthermore, this uniquely positions the Privileged-Plane Projective method as a tool for assessing whether neural refraction [2] occurs for out-of-training-distribution outlier representations.
@@
Using a comparative analysis between standard-Tanh and isotropic-Tanh, then similarly on Leaky-ReLU family of functions, this method can then determine whether there is a tendency for discrete representation clusters to form due to the task-agnostic functional form symmetries. It indicates that algebraic permutation equivariance tends to produce a quantising inductive bias on otherwise continuous representations, rather than task-necessitated representational collapses. As a consequence, it suggests that downstream phenomena, such as interpretable neurons, are not fundamental but rather a direct consequence of anisotropic functional forms. This may have benefits in AI safety and interpretable networks; however, preliminary evidence also suggests that this task-agnostic collapse can be pathological, such as through reduced representational capacity.

### v2

@@ Abstract
This paper presents a novel methodology for determining representational alignment, which builds upon the existing Spotlight Resonance method. Particularly, this new tool is used to gain insight into how discrete representations can emerge and organise in autoencoder models, through a controlled ablation study in which only the activation function is altered. Using this technique, the validity of whether function-driven symmetries can act as implicit inductive biases on representations is determined. Representations are found to tend to discretise when the activation functions are defined through a discrete algebraic permutation-equivariant symmetry. In contrast, they remain continuous under a continuous algebraic orthogonal-equivariant definition. This confirms the hypothesis: algebraic symmetries of network primitives can carry unintended inductive biases which produce task-independent artefactual structures in representations. The discrete symmetry of contemporary forms is shown to be a strong predictor for the induction of discrete representations transformed from otherwise continuous structures — a quantisation effect. This motivates further reassessment of functional forms in common usage. Moreover, this supports a general causal model for one mode in which discrete representations may form, and could constitute a prerequisite for downstream interpretability phenomena, including grandmother neurons, discrete coding schemes, general linear features and possibly Superposition. Hence, this tool and proposed mechanism for the influence of functional form on representations may provide insights into emergent interpretability research. Finally, preliminary results indicate that quantisation of representations appears to correlate with a measurable increase in reconstruction error, reinforcing previous conjectures that this collapse can be detrimental.
@@ 1.1 How Functional Form Choices May Result in Representational Inductive Biases
Many of the foundational functions in contemporary deep learning produce maps that differ as the direction changes; a property termed as anisotropy. This angular variation of the functions is hypothesised to make specific directions distinct, which the network then alters its representations about through optimisation. Hence, an angularly varying map could be expected to induce shaped angular representations. It is this connection between functional forms and representations which shall be explored in this work.
Directions singled out by a functional form’s structure may be termed a ‘privileged basis’ Elhage et al. (2022) or a set of ‘distinguished directions’ Bird (2025a). Such distinguished directions are commonplace through activation functions, normalisers, initialisers, regularisers, optimisers, architectures, operations, and more Bird (2025a). Hence, there are several avenues hypothesised through which representational influence may occur: directly through functional forms, indirectly through optimiser forms, and inherently through neuron connectivities. This proposed triad of influences would suggest that the representational alignment phenomenon may be systematic to deep learning’s anisotropic primitives and general practice. As a starting point, this work will isolate and determine the direct influence of activation functions as one of these causal influences on representations.
In particular, contemporary functional forms tend to make distinct the standard basis’ vectors, $\hat{e}_{i}$, and in such a way that each basis vector’s distinction is equal. This particular form of anisotropy can be categorised through their shared, underlying permutation symmetry construction Godfrey et al. (2023); Bird (2025a). This can now act as a defining characteristic relation for most contemporary functional forms. This is often due to their elementwise application, but appears more generally. The specific relation depends on the primitive in question; for most activation functions, this is an algebraic permutation equivariance Godfrey et al. (2023); Bird (2025a), as displayed in Eqn. 1. The hypothesis is deepened by suggesting that the maximal symmetries to which functional forms belong are a strong predictor of representational structure.
@@
The Spotlight Resonance Method Bird (2025b) has already empirically demonstrated how such activation functions can produce a tendency for representational alignments about the distinguished directions. This was achieved through transforms, and altering the number of the functional form’s distinguished directions, with corresponding changes in the representations mirroring the transformed structure of the distinguished directions being observed. The conclusion was that functional forms stimulate this tendency towards representational alignment, producing emergent task-agnostic alignment structure in representations, and hence establishing a causal relationship between the two. The observed structure consisted of a tendency for a greater density of representations to cluster about the functional form’s distinguished basis. This work also indicated the presence of ’grandmother neurons’, indicating how many clusters represent a distinct semantic and also associated their alignment with functional form choices. However, it did not establish whether discretisation itself was a phenomenon of this choice of form.
@@ 1.2 Motivations for This Study
Isotropic Deep Learning Bird (2025a) takes the consequences of such observations a step further. The Spotlight Resonance Method determined that task-agnostic structure in representations is associated with a functional form’s distinguished directions. The Isotropic Deep Learning approach proposes that since this structure transforms with the functional form, the anisotropic functional form choices may be the fundamental causal mode that induces the array of interpretable structures to begin with. In effect, anisotropic primitives provide an absolute frame, for each representation space, about which the representations organise and discretise into distinct clusters. Hence, questioning whether the ubiquity of anisotropy in contemporary deep learning constitutes an underlying origin for many such interpretable and emergent-structure phenomena, which are commonly observed — and whether this is beneficial or not. Furthermore, determining the validity of such a hypothesis would indicate whether these phenomena are truly fundamental to deep learning or stimulated by, and contingent on, the specific choices in functional form. This could have significant implications for the approach of network interpretability. Confirmation of this hypothesis would also demonstrate that such choices do carry inherent and unappreciated biases with their implementation.
This position motivates the comparative study of this work. Determining whether one can control and demonstrate the emergence and elimination of structure through broadening deep learning’s primitives to other form-defining choices, each providing its own inductive bias. One such proposed choice Bird (2025a) is by redefining primitives through an orthogonal algebraic-equivariance, displayed in Eqn. 2. This is termed an ‘isotropic’ definition of forms, and is a supergroup-symmetry of the original permutation symmetry, $S_{n}\subset\mathrm{O}\left(n\right)$. This defining choice of isotropic primitives is interesting, since it can eliminate all form-distinguished directions in a network. Therefore, functions developed to adhere to this modified symmetry definition can be used as a comparative mode to determine whether functional forms can induce these interpretable structures and indicate whether downstream alignment phenomena are attributable to anisotropic choices. If this is the case, it would demonstrate that these differing forms do carry unintended inductive biases, which would warrant major reconsideration of the forms used in standard practice.
$\forall\mathbf{R}\in\mathrm{O}\left(n\right),\forall\vec{x}\in\mathbb{R}^{n}\quad\mathbf{f}\left(\mathbf{R}\vec{x}\right)=\mathbf{R}\mathbf{f}\left(\vec{x}\right)$ (2)
@@
In particular, anisotropic forms may lead to a tendency for representations to overly cluster around distinguished directions. This is due to the anisotropic functional form’s nature, where some angular arrangements may become overrepresented through optimisation due to their irregularity. This task-agnostic over-density bias may be detrimental, as representations may tend to collapse towards single directions, such as aligned or anti-aligned, and consequently lose informative representational degrees of freedom, potentially damaging a network’s internal expressiveness of otherwise continuous structure. Hence, anisotropy may result in a quantising effect on representations, and thus also the semantics being represented — resulting in a limited representational capacity and neural coding scheme. Consequently, isotropic primitives are expected to remove this form-induced task-agnostic discretising bias on representations.
This may occur through several modes, hypothesised to stem from the underlying symmetry. To begin with, these symmetries are able to interact with the network structure and therefore its dynamics. This differs from symmetries in end-to-end networks, where their representations are designed to transform with the symmetry, resulting in the entire model being either equivariant or invariant to its effect. This is not the case in general networks, where a symmetry action does not pass cleanly through or identically cancel with layers. In general networks, a layer may ‘absorb’ a symmetry action in a non-invariant/equivariant manner, or have an interaction with its differing symmetry, meaning the symmetry can, in principle, ‘act’ on the network and hence produce a realisable effect. This is argued to result in various inductive biases that are considered underappreciated in significance. For example, the permutation-like symmetries about the standard basis result in an orthant partitioning of the space. Additionally, depending on the particular permutation symmetry, sets of these orthants may be analytically degenerate under hyperoctahedral transforms. The network’s optimisation will then shape representations about these orthants of differing and equal mapping, to achieve the desired computation. This results in representations that organise differently due to maps defined by the symmetry of the form, and in their extreme, may cluster over particular regions, producing a quantisation effect on representations. This is considered a general mode through which function-driven symmetries may influence representations as optimisation responds to their structure. The general approach is to study how various function-driven symmetries interact and act on the network and its representations, determining the various biases induced.
This mechanism would aid in explaining why the SRM study Bird (2025b) detected axis-aligned structure, since incidental alignment with directions people choose to define and observe as ‘individual neurons’ would have a vanishingly small likelihood of occurring in high-dimensional spaces. This is because there are vastly more orientations that do not align with the standard basis than those that do in high-dimensional spaces. However, if the analytic properties of permutation-based functions produce maps which differ along these directions, then the symmetry breaking of the space would single out such directions and make substantial alignment less peculiar. Therefore, it indicates an unintended bias of the model. A comparison between anisotropic and isotropic forms can directly probe such questions by adding and removing any distinction between directions.
Quantisation is argued to have potential pathological effects on the network. One heuristic is that the underlying distributions might be better represented continuously, as this aligns with the frequently continuous real-world semantics being represented, such as broad object morphologies and arangments, multivariate signals, lighting and shades, continuous expressions, modelling distributions, etc [^1] This aspect independently converges on similar motivation to Sabour et al. (2017)’s vector-valued routed capsule models but significantly diverges in terms of lack of capsules and routing and it is foremost concerning general architectures (potentially under reselection). Substantial differences arise also in the reconsideration of, and interactions between, sets of primitices, additionally the generality of the scope of such functional forms outside of just normalisation, alongside the investigation and leveraging of implicit inductive biases regarding various symmetries constructions. Therefore, this independent approach constitutes a considerably broader consideration of the impacts and generality of such forms and their symmetry-defining approach in deep learning Bird (2025a).. Therefore, enabling latent variables and representations to be continuous and potentially more interpolatable at various levels through a network, instead of an imposed task-agnostic structure, may be beneficial for many tasks. This may increase representational capacity and organisation of the representations.
@@
This connects to the coding scheme of the network. Axis-aligned representations, previously observed to frequently emerge Bird (2025b), approximate a local coding scheme for the network. This is where one neuron corresponds to one semantic. This coding scheme has decreased concept-interference, but has a much lower representational capacity than other coding schemes. Models have been demonstrated to produce other interference-capacity balances with various input sparsities Elhage et al. (2022); Bird (2025b), suggesting that axis-alignment is not an appropriate universal inductive bias — further motivating the study of alternative functional form inductive biases and their effects on representations.
@@
A functional form that does not distinguish particular directions may therefore not deter such continuous and interpolatable structures present in the dataset, and their preservation may be advantageous to the network. Yet, sometimes discretisation may be appropriate, such as in classification networks, where abstracting away redundant detail can be beneficial — though in autoencoding models this would not be an expectation. In any case, rigidly imposing this through a task-agnostic, function-driven, inductive bias about a specific arrangement might be detrimental. For example, when reducing the dimensionality of a dataset inherently containing discrete representations via a linear layer, to some arbitrary lower dimension, one would not expect it always appropriate for those discrete directions to be best represented through an alignment to an orthonormal basis or other imposed geometry for that space — as these are as arbitrary to the bottleneck chosen.
@@
Other speculative pathological modes are attributed to the hypothesised functional form quantisation. For example, discrete representations may be particularly susceptible to perturbations which move activations outside of the cluster, into an in-between region of representation space. Maps from these unpopulated regions may be poorly trained and fragile, due to their training-time rarity, and therefore highly unpredictable in their map. This may suggest that such a quantising inductive bias decreases adversarial robustness, whereas removing this bias through isotropy may be advantageous. However, this speculation is not explored within this work.
This paper aims to assess the validity of the former prediction. The motivating questions are: do discrete-symmetry defined functional forms tend to produce approximately discrete representations, quantising an otherwise underlying continuous data structure, whilst continuous-symmetry defined functional forms may tend to preserve a more continuous representation. Furthermore, is the prior construction potentially pathological in nature? Do continuous representations tend to improve performance, since quantised representations may harm expressiveness? Therefore, determining the validity of this prediction may provide crucial insight into the potential of Isotropic deep learning and the emergence of downstream interpretability structure, namely alignment, grandmother neurons, general interpretable directions, discrete codes, and possibly Superposition Elhage et al. (2022) and Neural Collapse Papyan et al. (2020).
@@
Most importantly, this study aims to clarify the link between functional form algebraic symmetries and their impact on representations. It will first be demonstrated that symmetry is a key predictor of the emergent task-agnostic structure in representations. This encourages an exploration into the effects of various symmetry groups and possibly a reconsideration of functional form primitives, network-wide, if alternative symmetries are beneficial. This is because symmetry can act as a classifying and definitional approach to functional forms.
@@ 1.3 Methodological Approach
Instead of rotating the privileged basis or altering the number of distinguished directions to determine alignment of representations, this study will compare representations between the presence and absence of any distinguished directions to determine if this is a causal mechanism for the production of approximately discrete representations. Due to the expected systematic nature of this inductive bias, care was taken to prevent the unintentional reintroduction of anisotropy. To achieve this, distinguished directions will be removed network-wide to avoid these confounding and obscuring interactions. This is achieved by utilising isotropic primitives and an autoencoder model for reconstruction  [^2] The use of an autoencoder model prevents anisotropy through choices in the output layer representation. For example, one-hot encoding in classification outputs is also a choice in form defined through permutation symmetry; therefore, it could induce further spurious structure. A reconstruction task using autoencoders may therefore mitigate this and enable more data-representative and ‘natural’ embeddings. Moreover, it has been speculated Bird (2025b, a) whether the neural-collapse classification phenomenon Papyan et al. (2020) is related to the proposed mechanisms. So autoencoder reconstruction provided a suitable minimalistic setting, free from these influences to study form-induced representational quantisation., except for an activation function which can selectively reintroduce anisotropy through changes to its form. Hence, this isolates the causality of the phenomenon to the activation functions.
This requires two activation functions, which are highly similar in action but differ in the presence or absence of anisotropy. The isotropic position paper Bird (2025a) provides such a preliminary isotropic activation function, which enables this comparative analysis. An analogue of standard elementwise (anisotropic) tanh, displayed in Eqn. 3, can be compared with its orthogonal analogue Isotropic-Tanh, displayed in Eqn. 4 — each is displayed in multivariate form for comparison $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. These functions produce equivalent non-linear maps along the standard basis directions: $f_{\text{iso}}\left(\alpha\hat{e}_{i}\right)=f_{\text{aniso}}\left(\alpha\hat{e}_{i}\right)$ forall $\hat{e}_{i}$ and $\alpha$. Standard-Tanh is algebraically equivariant to the hyperoctahedral group (standard-basis permutation with sign-flips), $B_{n}$, whilst isotropic-tanh is algebraically equivariant to the orthogonal group, $\mathrm{O}\left(n\right)$, with $B_{n}\subset\mathrm{O}\left(n\right)$.
@@
This approach is also extended to a Leaky-ReLU Maas et al. (2013) family, which are modified to adhere to the following symmetries $S_{n}\subset\mathrm{O}\left(n\right)$. This is discussed further in App. B.4 and indicates that algebraic symmetries are the primary predictor, since all other algebraic properties of Leaky-ReLU substantially differ from the tanh-like family of Eqns. 3 and 4.
This work also modifies the prior Spotlight Resonance Method (SRM). This modified tool is termed the Privileged-Plane Projective Method (PPP method), which operates on a similar principle. However, unlike SRM, this adaptation does not project out the vector length in the process. Hence, information regarding both angular and magnitude distributions is retained. This provides a more nuanced understanding of the network’s representational structures; however, SRM may still be preferable in determining statistical measures of the angular representations. Furthermore, this uniquely positions the Privileged-Plane Projective method as a tool for assessing whether neural refraction Bird (2025a) occurs for out-of-training-distribution outlier representations.
@@
Using a comparative analysis between standard-Tanh and isotropic-Tanh, and similarly on Leaky-ReLU family of functions, this method can then determine whether there is a tendency for discrete representation clusters to form due to the task-agnostic functional form symmetries. It indicates that algebraic permutation equivariance tends to produce a quantising inductive bias on otherwise continuous representations, rather than task-necessitated representational collapses. As a consequence, it suggests that downstream phenomena, such as interpretable neurons, are not fundamental but rather a direct consequence of anisotropic functional forms. This may have benefits in AI safety and interpretable networks; however, preliminary evidence also suggests that this task-agnostic collapse can be pathological, such as through reduced representational capacity.